OpenStudy (anonymous):

11. You are given a block that has a mass of 3.42 grams. You measure it’s volume to be 2.00 cm by 5.00 cm by 4.00 cm. Calculate the density of the block.

OpenStudy (jdoe0001):

density will be \(\bf \cfrac{weight}{volume} \implies \cfrac{3.42gms}{2cm\times 5cm \times 4cm}\)

OpenStudy (anonymous):

ok solve that equation thing right ok thank you

OpenStudy (anonymous):

so 0.0855 would be correct, checking with u.

OpenStudy (anonymous):

what would be the unit?

OpenStudy (jdoe0001):

well, what do you think will be the unit? notice the fractions, they have their units :)

OpenStudy (anonymous):

grams? i believe

OpenStudy (jdoe0001):

grams over ?

OpenStudy (anonymous):

over? lol cenimeters?

OpenStudy (anonymous):

or do u mean grams squared?

OpenStudy (anonymous):

lol tell me this is for test review that i already finished the test, so its worthless, but its required to turn in, i got 35/35 on my test btw

OpenStudy (jdoe0001):

\(\bf \cfrac{weight}{volume} \implies \cfrac{3.42gms}{2cm\times 5cm \times 4cm} \implies \large 0.0855 \frac{grams}{cm^3}\) so, that means, that, there are 0.0855 grams per each cubic centimeter

OpenStudy (anonymous):

oh ok thank you :D

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

gms^3 is that right? lol

OpenStudy (jdoe0001):

\(\bf cm \times cm \times cm \implies cm^3\)

OpenStudy (anonymous):

so no grams, just cm^3 so 0.0855cm^3?

OpenStudy (jdoe0001):

\(\bf \cfrac{weight}{volume} \implies \cfrac{3.42gms}{2cm\times 5cm \times 4cm} \implies \large 0.0855 \color{red}{\frac{grams}{cm^3}}\)

OpenStudy (anonymous):

ok gotcha, thanks for the help :D

OpenStudy (jdoe0001):

np